The distance from A to B is 450km. If the journey from A to B takes 9 hours by bus but only 5 hours by car, how many kilometers per hour does the car travel faster than the bus?
Answer by Theo(13342)

rate * time = distance.
let b = the rate of the bus
let c = the rate of the car
b * 9 = 450 leads to:
b = 450 / 9 = 50 kmph for the bus
c * 5 = 450 leads to:
c = 450 / 5 = 90 kmph for the car
the car is traveling 90 - 50 = 40 kmph faster than the bus.